The radius and surface charge density of a uniformly charged spherical shell are 20 cm and ,m C0.3 —2 u respectively. Calculate the electric potential at a distance (a) 40 cm and (b) 15 cm from the center of the shell. 



1
Expert's answer
2021-06-01T10:47:47-0400
"q=4\pi (0.2)^2(0.3\cdot10^{-6})=1.5\cdot10^{-7}C"

a)


"V=9\cdot10^9\frac{1.5\cdot10^{-7}}{0.4}=3375\ V"

b)


"V=9\cdot10^9\frac{1.5\cdot10^{-7}}{0.2}=6750\ V"